# MERIT: Multilingual Semantic Retrieval with Interleaved Multi-Condition Query

### v2

@@ B.3 Quality Control and Validation Protocol
  • (3) Query Formulation Compliance Protocol. To ensure statistical robustness and ecological validity, every search query must be generated through our tri-modal sampling methodology: (i) Conventional Uniform Sampling, (ii) Attribute Uniform Sampling, and (iii) High-similarity Product Prioritization Sampling, as detailed in Appendices A.3.1-A.3.3. Cold-start expansion procedures (Appendix A.3.4) shall be employed only after primary sampling methods have been exhausted. Each query composition undergoes verification against pre-established diversity metrics.
  • (4) Multi-stage Validation Protocol. All dataset entries are subject to mandatory dual-phase validation: (i) Automated Validation Phase—employing deterministic rule-based systems and statistical outlier detection methods to identify and eliminate non-conforming samples; followed by (ii) Expert Validation Phase—requiring assessment by annotators with demonstrated proficiency in the relevant languages and product domains. Samples must achieve unanimous approval through both validation phases to MERIT inclusion in the final dataset. Non-compliant samples shall be documented for methodological refinement purposes.
B.4 Handling Ambiguities
@@ B.5 Ethical Considerations
B.6 Data Contamination Considerations
@@ B.7 Annotation Platform
B.8 Data Preparation and Release
@@
Appendix C More Dataset Analysis
C.1 Global Statics
@@ Appendix D Dataset Examples
Appendix E Experiments Details
E.1 Baselines
@@
E.2 Other Dataset Usage
@@
E.3 Main Experiments Settings
@@
E.4 More Experiments Results
@@
Out of Distribution Scenarios. As mentioned in Sec. 3.3, we tested several out-of-distribution (OOD) scenarios. Tables 87, and 9 correspond to the Zero-shot, OOD, and Mixed results depicted in Figure 6(b), respectively. Zero-shot refers to direct inference using Qwen2.5-VL’s [EOS] token. OOD indicates that for each row in Table 7, we excluded the specified OOD data (e.g., for the first row concerning Language ID OOD, we removed all queries containing the ID language from the training set, while for the test set, we only selected data where all languages were ID). Mixed refers to training conducted on the complete training dataset.
@@
E.5 Error Analysis Details
@@ E.6 Human preference
E.7 Condition Count on MERIT
As demonstrated in Table 12, we observe that the performance of nearly all models degrades with increasing condition count. Notably, under the most complex scenario with $4$ conditions, even the best-performing model achieves only $36.84$ R@5, highlighting the inherent difficulty posed by the combinatorial complexity of multi-condition queries.
@@ E.7 Condition Count on MERIT
Appendix F Broader Impact
F.1 Impact

### v3

@@ 7.3 Quality Control and Validation Protocol
  • (3) Query Formulation Compliance Protocol. To ensure statistical robustness and ecological validity, every search query must be generated through our tri-modal sampling methodology: (i) Conventional Uniform Sampling, (ii) Attribute Uniform Sampling, and (iii) High-similarity Product Prioritization Sampling, as detailed in Appendices 6.3.1-6.3.3. Cold-start expansion procedures (Appendix 6.3.4) shall be employed only after primary sampling methods have been exhausted. Each query composition undergoes verification against pre-established diversity metrics.
  • (4) Multi-stage Validation Protocol. All dataset entries are subject to mandatory dual-phase validation: (i) Automated Validation Phase – employing deterministic rule-based systems and statistical outlier detection methods to identify and eliminate non-conforming samples; followed by (ii) Expert Validation Phase – requiring assessment by annotators with demonstrated proficiency in the relevant languages and product domains. Samples must achieve unanimous approval through both validation phases to MERIT inclusion in the final dataset. Non-compliant samples shall be documented for methodological refinement purposes.
7.4 Handling Ambiguities
@@ 7.5 Ethical Considerations
7.6 Data Contamination Considerations
@@ 7.7 Annotation Platform
7.8 Data Preparation and Release
@@
8 More Dataset Analysis
8.1 Global Statics
@@ 9 Dataset Examples
10 Experiments Details
10.1 Baselines
@@
10.2 Other Dataset Usage
@@
10.3 Main Experiments Settings
@@
10.4 More Experiments Results
@@
Out of Distribution Scenarios. As mentioned in Sec. 3.3, we tested several out-of-distribution (OOD) scenarios. Tables 8, 7, and 9 correspond to the Zero-shot, OOD, and Mixed results depicted in Figure 6(b), respectively. Zero-shot refers to direct inference using Qwen2.5-VL’s [EOS] token. OOD indicates that for each row in Table 7, we excluded the specified OOD data (e.g., for the first row concerning Language ID OOD, we removed all queries containing the ID language from the training set, while for the test set, we only selected data where all languages were ID). Mixed refers to training conducted on the complete training dataset.
@@
10.5 Error Analysis Details
@@ 10.6 Human preference
10.7 Condition Count on MERIT
As demonstrated in Table 12, we observe that the performance of nearly all models degrades with increasing condition count. Notably, under the most complex scenario with four conditions, even the best-performing model achieves only $36.84$ R@5, highlighting the inherent difficulty posed by the combinatorial complexity of multi-condition queries.
@@ 10.7 Condition Count on MERIT
11 Broader Impact
11.1 Impact